A random sample is size n = 2 are drawn from a finite population consisting of the numbers 3,4,5,6 and 7


1
Expert's answer
2021-04-15T06:30:42-0400

1. Mean


"\mu=\dfrac{3+4+5+6+7}{5}=5"

Variance


"\sigma^2=\dfrac{1}{5}\big((3-5)^2+(4-5)^2+(5-5)^2"

"(6-5)^2+(7-5)^2\big)=2"


Standard deviation


"\sigma=\sqrt{\sigma^2}=\sqrt{2}\approx1.4142"



2. We have population values "3,4,5,6,7," population sizeΒ "N=5" and sample sizeΒ "n=2." Thus, the number of possible samples which can be drawn without replacement is


"\dbinom{N}{n}=\dbinom{5}{2}=10""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 3,4 & 3.5 \\
 \hdashline
 2 & 3,5 & 4 \\
 \hdashline
 3 & 3,6 & 4.5 \\
 \hdashline
 4 & 3,7 & 5 \\
 \hdashline
 5 & 4,5 & 4.5 \\
 \hdashline
 6 & 4,6 & 5 \\
 \hdashline
 7 & 4,7 & 5.5 \\
 \hdashline
 8 & 5,6 & 5.5 \\
 \hdashline
 9 & 5,7 & 6 \\
 \hdashline
 10 & 6,7 & 6.5 \\ \hline
\end{array}"

3. The sampling distribution of the sample means.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 7/2 & 1& 1/10 & 7/20 & 49/40 \\
 \hdashline
 4 & 1 & 1/10 & 8/20 & 64/40 \\
 \hdashline
 9/2 & 2 & 2/10 & 18/20 & 162/40 \\
 \hdashline
 5 & 2 & 2/10 & 20/20 & 200/40 \\
 \hdashline
 11/2 & 2 & 2/10 & 22/20 & 242/40 \\
 \hdashline
 6 & 1 & 1/10 & 12/20 & 144/40 \\
 \hdashline
 13/2 & 1& 1/10 & 13/20 & 169/40 \\
 \hdashline
 
 Total & 10 & 1 & 100/20 & 1030/40 \\ \hline
\end{array}"

4.


"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{100}{20}=5"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.


"E(\bar{X})=5=\mu"

5.

"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"

"=\dfrac{1030}{40}-(\dfrac{100}{20})^2=\dfrac{3}{4}=0.75"

"\sqrt{Var(\bar{X})}=\sqrt{\dfrac{3}{4}}=\dfrac{\sqrt{3}}{2}\approx0.8660"

Verification:

"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{2}{2}(\dfrac{5-2}{5-1})"

"=\dfrac{3}{4}=0.75, True"